1. Fisher Group offers credit to its customers and charges interest of 2 percent per month. What is the annual percentage rate?

2. A bond currently has a yield to maturity of 8 percent. The bond matures in 3 years and pays interest semi-annually. The coupon rate is 7 percent. What is the current price of this bond?
Answer #1

1. annual percentage rate = (interest rate/no. of days)*365 = (0.02/30)*365 = 6.6666666666666666666666666666667e-4*365 = 0.2433 or 24.33% rounded off to 24%

30 is no. of days in a month and 365 is the no. of days in year.

2. we can use financial calculator for calculation of price of the bond. let's assume face value of the bond is $1,000.

due to interest paid semi-annually, maturity will be doubled and yield to maturity & coupon will be half.

N = maturity = 3*2 = 6; I/Y = yield to maturity = 8%/2 = 4%; PMT = coupon = $1,000*7%/2 = $35; FV = face value = $1,000 > CPT = compute > PV = current price = $973.79

the current price of this bond is $973.79.